# Towards Robust Stability Prediction in Smart Grids: GAN-based Approach under Data Constraints and Adversarial Challenges

### v1

@@ Abstract
Smart grids are critical for addressing the growing energy demand due to global population growth and urbanization. They enhance efficiency, reliability, and sustainability by integrating renewable energy. Ensuring their availability and safety requires advanced operational control and safety measures. Researchers employ AI and machine learning to assess grid stability, but challenges like the lack of datasets and cybersecurity threats, including adversarial attacks, persist. In particular, data scarcity is a key issue: obtaining grid instability instances is tough due to the need for significant expertise, resources, and time. However, they are essential to test novel research advancements and security mitigations. In this paper, we introduce a novel framework to detect instability in smart grids by employing only stable data. It relies on a Generative Adversarial Network (GAN) where the generator is trained to create instability data that are used along with stable data to train the discriminator. Moreover, we include a new adversarial training layer to improve robustness against adversarial attacks. Our solution, tested on a dataset composed of real-world stable and unstable samples, achieve accuracy up to 97.5% in predicting grid stability and up to 98.9% in detecting adversarial attacks. Moreover, we implemented our model in a single-board computer demonstrating efficient real-time decision-making with an average response time of less than 7ms. Our solution improves prediction accuracy and resilience while addressing data scarcity in smart grid management.
Index Terms:
@@
Smart Grid, Stability Prediction, Generative Adversarial Networks (GANs), Adversarial Attacks.
I Introduction
@@
The rapid global population growth, economic expansion, and urbanization are projected to significantly increase energy consumption, coinciding with the adoption of renewable energy. Traditionally, energy grids functioned with a unidirectional flow from production centers to consumers. However, the emergence of the so-called prosumers —entities that both consume and supply energy—requires a shift to bidirectional energy flow within grids [1]. In this context, smart grids represent a transformative solution, enhancing the efficiency, reliability, and sustainability of electrical grids through advanced technologies. They create a modern electricity distribution network characterized by improved dependability, effectiveness, sustainability, and bidirectional communication capabilities [2]. Smart grids leverage real-time data, communication networks, and intelligent control mechanisms to optimize electricity generation, distribution, and consumption, enabling a responsive energy ecosystem with a two-way flow of information between utilities and customers [2]. As the energy landscape evolves, smart grids are vital for integrating renewable resources like solar and wind, mitigating variability, and promoting overall sustainability. Their adaptability is crucial for incorporating new technologies and managing operational characteristics, such as power collection timing and delivery capacity [3].
@@
Accurately predicting renewable energy generation is critical for maintaining stable and efficient power system operations, especially when managing the inherent fluctuations of sustainable energy sources [4]. Similarly, robust forecasting techniques are essential to preempt disruptions in balancing electricity supply and demand. To address the challenges posed by fluctuating power grids, various smart grid strategies have been developed, with a key focus on balancing supply and demand. One of the central concepts in these strategies is the regulation of consumer demand, commonly referred to as the demand response strategy [5, 6]. Demand response involves adjustments in electricity consumption by customers in response to changes in electricity prices, diverging from their typical consumption patterns [7]. A promising innovation in this domain is Decentralized Smart Grid Control (DSGC), which integrates electricity prices with grid frequency, a parameter that can be measured by prosumers [8]. In this framework, frequency increases during power surpluses and decreases during shortages [9], enabling real-time pricing mechanisms that allow prosumers to dynamically adjust their demand. However, the effective implementation of DSGC requires overcoming several challenges, including maintaining grid stability amidst rapid price fluctuations, addressing varying price sensitivities among participants, and accommodating differences in their response times [10].
@@
Grid instability can cause major disruptions to electricity supply, impacting everyday activities and economic systems. When the power grid experiences instability, it may lead to outages, harm electrical equipment, and create safety risks. For example, variations in voltage can make lights flicker and potentially harm delicate devices [11]. In extreme situations, a localized outage caused by its instability, within the grid can trigger a cascading effect, resulting in extensive blackouts [3]. In the real world, a fault in an underground cable caused a near-total blackout in Puerto Rico, leaving approximately 1.3 million people without electricity [12].
@@
Machine Learning (ML) and Artificial Intelligence (AI) have proven to be highly effective in stability prediction in decentralized smart grids, where various models have been employed achieving accuracies close to 100% in detecting unstable samples [1, 13, 14, 15]. However, significant challenges persist, particularly the absence of datasets containing both stable and unstable instances. This gap poses a major obstacle to advancing the field with AI, as such datasets are crucial for effective model training. This scarcity arises from the undesirable and rare nature of instability, which typically indicates system failures, malfunctions, and fluctuations in energy demand. Collecting real-world data is challenging and risky, as deliberately inducing instability for data collection or model training could lead to disruptions, equipment damage, or safety hazards. Additionally, acquiring sufficient data is time-consuming since it could require human effort for labelling [16].
@@
As smart grids increasingly depend on data-driven technologies, robust security measures are crucial to protect confidentiality, integrity, and availability within the energy infrastructure [17]. AI and ML techniques offer advanced predictive capabilities and improve grid management efficiency, but their vulnerability to adversarial attacks remains a critical and often overlooked issue in the literature [18]. The extensive interconnection between devices and remote access points expands the attack surface, creating potential entry points for attackers to infiltrate the entire network. Compromising a single node can enable attackers to inject malicious data into the smart grid data flow, highlighting the need for enhanced security protocols. Several studies [19, 20] are investigating the susceptibility of AI-enabled stability prediction systems in smart grids to adversarial attacks. However, no one is addressing stability prediction and resistance to adversarial attacks using only one model.
@@
Contributions. In this paper, we propose GAN-Stability, a novel framework for stability prediction and adversarial attack detection. For training, our solution uses only one class (stable samples) from a two-class real dataset of the stability prediction task. Our solution trains the discriminator of a Generative Adversarial Network (GAN) with i) synthetic samples generated by the GAN generator that are potentially out-of-distribution relative to the stable class, ii) stable samples from a real-world dataset; and iii) adversarial samples made by attacking stable data. Thanks to the synthetic generation of potential unstable samples, we alleviate the need to render the smart grid unstable to collect data. We also reduce the need for an additional model to handle adversarial attacks by integrating adversarial sample detection directly into the GAN framework. Our approach trains the discriminator to identify adversarial samples as part of the unstable class within the same process, simplifying the overall architecture. We tested our framework over a real-world dataset containing both stable and unstable data. Our solution achieves an accuracy of up to 0.975 in stability prediction despite the absence of instability instances in the training data. We also assess the robustness of our adversarial training approach against state-of-the-art attacks in whitebox and greybox scenarios, achieving up to 0.989 accuracies in classifying attacks as instability conditions. While reducing the burden on dataset development, our solution eliminates the need for a separate Anomaly Detection System (ADS) to detect adversarial attacks.
@@
  • We design GAN-Stability, a framework to train stability prediction systems employing exclusively stable instances while retaining high accuracies. To the best of our knowledge, we are the first to develop such a model in the smart grid context.
  • We enhanced the capabilities of our model with adversarial training, providing GAN-Stability with the capabilities to classify adversarial attacks as unstable samples without the need for an external ADS.
  • We evaluate our system on a widely used real-world dataset. We show an accuracy of up to 0.975 for our stability prediction model to identify unstable conditions. Furthermore, the addition of adversarial training allows the model to identify state-of-the-art adversarial attacks with an accuracy of up to 0.989.
  • We compare our solution with state-of-the-art models that require both stable and unstable data for training. Our model shows increased or comparable results despite the reduced data requirement for training (i.e., only stable data are needed).
  • We test GAN-Stability on a cheap microcomputer, showing acceptable time requirements for training and an average response time of less than 7ms for testing.
  • We make the code of our systems, attacks, and the dataset available at: https://anonymous.4open.science/r/GAN-Stability-10CE
Organization. The rest of the paper is organized as follows: Section II reviews existing stability prediction systems and their associated security implications. Section III introduces the system and threat model, reflecting real-world scenarios encountered during training and potential attacks. Section IV discusses the adversarial attack methodologies employed in this study. Section V details the proposed stability prediction system and its methodology, while Section VI evaluates our proposed system using different criteria. Finally, Section VII concludes the study.
II Related Works
In this section, we delve into the existing literature concerning stability prediction systems and their security implications. Specifically, we scrutinize established methodologies for stability prediction in Section II-A, while we assess the current landscape of attacks targeting these systems in Section II-B.
II-A Smart Grid Stability Prediction using AI
@@ II-B ML Adversarial Attacks
The paper by Tian et al. [31] investigates security concerns of neural network-based state estimation in smart grids, focusing on adversarial attacks and proposing an efficient method for executing these attacks. Sayghe et al. [32] investigate the impact of adversarial examples on the detection of False Data Injection Attacks (FDIAs) using DL algorithms. Their research examines the repercussions on MLP when exposed to two different adversarial attack strategies. Ahmadian et al. [33] introduced a FDIA using a GAN framework, where the attacker acts as the generative network and the Energy System Operator (ESO) serves as the discriminative network. The attacker generates deceptive data to evade detection by the power system state estimator through an optimization process. Li et al. [34] show that well-established ML models used in energy theft detection systems are susceptible to adversarial attacks. They develop a method to create adversarial measurements, allowing attackers to report significantly lower power consumption to utility companies and evade detection by ML-based systems. Despite various research works investigating adversarial attack effects in this context, no one ever considers mitigation techniques, especially merged with stability prediction.
III System and Threat Model
@@ III-A System Model
In an operational setting without active threats, a DSGC stability prediction system assesses whether the grid remains stable or unstable, particularly in a decentralized smart grid context where electricity prices are tied to grid frequency that carry all necessary information about the current power balance. In fact, the stability of electrical grids depends on the balance between electricity generation and demand [7, 10]. In the context of DSGC, stability is characterized by synchronized node frequencies ($\omega$) and steady power flows ($P_{jk}$) across the grid. Stability requires minimal angular frequency deviations and effective damping to suppress oscillations. Stability is evaluated by the system’s ability to return to equilibrium after disturbances, as measured by mathematical models such as local stability (i.e., linear stability exploring dynamical stability around the steady-state operation of the grid) [10]. Instability arises when synchronization is lost, resulting in significant frequency deviations, amplified oscillations, and destabilized power flows. These effects can lead to cascading failures, particularly when delays, resonance effects, or insufficient damping prevent the grid from recovering.
@@
Our system leverages ML and AI algorithms to perform binary classification to categorize grid samples into stable or unstable classes using various input data that the control center collects from nodes on the grid. Examples of such data include the reaction time of each participant, which indicates how quickly consumers or systems respond to changes; the price elasticity coefficients which reflect how sensitive power consumption is to changes in electricity prices; and nominal power consumption and production features which represent the baseline levels of power used or generated by the system. As we discuss in Section V-A, in our system we employ 12 different features which are usually collected by cheap equipment by a particular prosumer and sent to a control center. Before deployment, the model is trained on clean, uncorrupted data to ensure reliable predictions. Stable grid instances are easy to obtain from operational data since represent the vast majority of operation time, while gathering instability data is more difficult. Instability instances require careful labeling by human experts, and acquiring a comprehensive dataset involves long-term observation and considerable resources.
@@
In our system model, we adopt a pragmatic approach where we collect enough instances of stability from cheap equipment by particular prosumers [7]. These instances, all from the same stable label, serve as the sole data for training the stability prediction model. Therefore, to ensure the model’s effectiveness, we should collect a sufficient amount of data that is both comprehensive and representative of the underlying distribution of stable grid conditions. By focusing solely on stability instances, we streamline the training process and alleviate the need for extensive data collection efforts associated with capturing instances of instability. To the best of our knowledge, all the models in the existing literature typically rely on access to both labels in the dataset to make accurate predictions [1, 3, 14, 35, 36, 37]. Our approach challenges this conventional paradigm by demonstrating that accurate stability prediction can be achieved using only instances of stability for training.
III-B Threat Model
@@
  • White-box Scenario: In this scenario, the attacker possesses comprehensive access to both the data employed in testing the model and detailed information regarding the model’s architecture and parameters. This advantageous position provides the attacker with ample opportunities to exploit vulnerabilities in the system. By leveraging this intelligence, the attacker can meticulously craft powerful adversarial samples aimed at deceiving the model. Such a scenario could occur if an attacker compromises the control center of a distributed smart grid, for example, through malware infiltration or physical attacks targeting power companies [39].
@@
  • Grey-box Scenario: In this scenario, the attacker has access to the testing data but lacks access to the architecture and parameters of the main model. Despite this limitation, the attacker can still conduct evasion attacks by employing a surrogate model —an alternative model trained on the same dataset— with different architectures. The effectiveness of these attacks relies on transferability properties or the chosen architecture. This condition can exist if an attacker compromises enough prosumers or entities in the system, thus gaining access to several data but without knowledge of the actual employed model. IoT botnets are an example of how such scenarios could easily become a reality [40].
@@
    To generate adversarial samples, we utilize the LSTM model proposed in [3] as surrogate model. These adversarial samples will then be deployed against our primary stability prediction system (GAN-Stability). This setup simulates a real-world scenario where potential attackers have access to limited information about the system.
IV Reference Attacks

### v2

@@ Abstract
Smart grids are crucial for meeting rising energy demands driven by global population growth and urbanization. By integrating renewable energy sources, they enhance efficiency, reliability, and sustainability. However, ensuring their availability and security requires advanced operational control and safety measures. Although artificial intelligence and machine learning can help assess grid stability, challenges such as data scarcity and cybersecurity threats, particularly adversarial attacks, remain. Data scarcity is a major issue, as obtaining real-world instances of grid instability requires significant expertise, resources, and time. Yet, these instances are critical for testing new research advancements and security mitigations. This paper introduces a novel framework for detecting instability in smart grids using only stable data. It employs a Generative Adversarial Network (GAN) where the generator is designed not to produce near-realistic data but instead to generate Out-Of-Distribution (OOD) samples with respect to the stable class. These OOD samples represent unstable behavior, anomalies, or disturbances that deviate from the stable data distribution. By training exclusively on stable data and exposing the discriminator to OOD samples, our framework learns a robust decision boundary to distinguish stable conditions from any unstable behavior, without requiring unstable data during training. Furthermore, we incorporate an adversarial training layer to enhance resilience against attacks. Evaluated on a real-world dataset, our solution achieves up to 98.1% accuracy in predicting grid stability and 98.9% in detecting adversarial attacks. Implemented on a single-board computer, it enables real-time decision-making with an average response time of under 7ms.
keywords:
@@
Smart Grids, Adversarial Attacks , Stability Prediction, Generative Adversarial Networks (GANs).
journal: Internet of Things
[…]
{graphicalabstract}
{highlights}
Tackles smart grid stability detection while addressing data scarcity.
Proposes a novel GAN framework detects grid instability using only stable data.
Generator creates OOD samples beyond the stable class to train the discriminator.
Added adversarial training layer to classify attacks as instability instances.
Deployed on single-board computer with under 7 ms avg response time.
1 Introduction
@@
The rapid growth of the global population, economy, and urban areas is expected to significantly increase energy demand, coinciding with the rise of renewable energy. Traditionally, energy grids featured a unidirectional flow from producers to consumers. But the emergence of prosumers—entities that both consume and supply energy—demands a shift to bidirectional flow [1]. Smart grids offer a transformative solution, enhancing the efficiency, reliability, and sustainability of electrical networks through advanced technologies. They enable modern electricity distribution with improved dependability, effectiveness, and two-way communication [2]. As the energy landscape evolves, smart grids are key to integrating renewable energies like solar and wind, addressing variability, and supporting sustainability. Their adaptability is crucial for managing new technologies and operational features, such as power collection timing and delivery capacity [3].
@@
Accurate forecasting of renewable energy generation is vital for stable and efficient power system operations, particularly given the inherent variability of sustainable sources [4]. Similarly, robust forecasting methods help preempt disruptions in balancing electricity supply and demand. To tackle the challenges of fluctuating power grids, smart grid strategies have emerged, with a key focus on supply-demand balancing. A central approach is the demand response strategy, where consumers adjust electricity use in response to price changes, deviating from typical consumption patterns [5, 6, 7]. A promising advancement in this area is Decentralized Smart Grid Control (DSGC), which integrates electricity prices to grid frequency—a measurable parameter for prosumers [8]. Frequency rises during surpluses and falls during shortages [9], enabling real-time pricing that prompts prosumers to adapt demand dynamically. However, effective DSGC implementation faces several challenges, such as ensuring grid stability amid rapid price shifts, handling varied price sensitivities, and accommodating differences in participant response times [10].
@@
Grid instability can cause major disruptions to the electricity supply, affecting daily life and economic systems. When the power grid becomes unstable, it can lead to outages, damage electrical equipment, and pose safety risks. For example, voltage fluctuations may cause lights to flicker and potentially damage sensitive devices [11]. In extreme situations, a localized outage triggered by grid instability can cascade into widespread blackouts [3]. A real-world example is the near-total blackout in Puerto Rico, where a fault in an underground cable left approximately 1.3 million people without power [12].
@@
Machine Learning (ML) and Artificial Intelligence (AI) have proven highly effective for stability prediction in decentralized smart grids, with several models achieving near-perfect accuracy in detecting unstable samples [1, 13, 14, 15]. However, a major challenge remains: the lack of real-world datasets containing both stable and unstable instances. This gap poses a major obstacle to advancing the field with AI, as such datasets are essential for training reliable models. The scarcity arises from the fact that unstable behaviors—typically indicating system failures, malfunctions, or sudden demand fluctuations—are both rare and undesirable. Collecting such data in a real-world scenario is difficult and risky, as inducing unstable behaviors could cause serious disruptions, equipment damage, or safety issues. Moreover, building high-quality datasets is time-consuming and often requires manual labeling [16].
@@
Beyond data scarcity, another major challenge in smart grid applications is ensuring the security of AI-driven systems. As smart grids increasingly rely on data-driven technologies, robust security measures are crucial to protect the confidentiality, integrity, and availability of energy infrastructure [17]. While AI and ML techniques enhance predictive capabilities and grid management efficiency, their vulnerability to adversarial attacks remains a critical and often overlooked issue in the literature [18]. The extensive interconnection between devices and remote access points expands the attack surface, creating potential entry points for attackers to infiltrate the entire network. Several studies [19, 20] have examined the susceptibility of AI-based stability prediction systems in smart grids to such attacks. For example, Efatinasab et al. [3] demonstrated how an attacker can exploit grid stability prediction models by injecting adversarial data, causing the system to misclassify unstable conditions as stable. This threat is particularly severe during high-demand periods, such as extreme weather events [21], when power systems are under stress and require corrective measures. If misleading data prevents appropriate responses—like load shedding or activating backup systems—overloads and failures in critical components may occur. However, to the best of our knowledge, no existing work addresses stability prediction using only stable data while also ensuring robustness against adversarial attacks within a single integrated model.
@@
Contributions. In this paper, we propose GAN-Stability, a novel framework for stability prediction and adversarial attack detection. Our approach is trained using only one class (stable samples) from a two-class dataset for the stability prediction task. Our solution trains the discriminator of a Generative Adversarial Network (GAN) with three types of data: i) synthetic samples generated by the GAN generator that are potentially Out-Of-Distribution (OOD) relative to the stable class, ii) stable samples from a real-world dataset, and iii) adversarial samples crafted by attacking the stable data. By generating OOD samples synthetically, our method alleviates the need to induce unstable behaviors in the smart grid for data collection. These generated samples may reflect unstable behavior, measurement anomalies, fault-induced disturbances, or other deviations from normal behavior. Moreover, GAN-Stability integrates adversarial detection directly into the GAN framework, removing the need for a separate model. Our approach trains the discriminator to identify adversarial samples as part of the unstable class within the same process, simplifying the overall architecture and improving robustness. We tested our framework on a dataset containing both stable and unstable data. Our solution achieves an accuracy of up to 0.981 in stability prediction, even when trained exclusively on stable data. We also evaluate the robustness of our adversarial training approach against state-of-the-art attacks in both whitebox and greybox scenarios, achieving accuracies up to 0.989 in classifying attacks as unstable behaviors. While reducing the burden on dataset development, our solution eliminates the need for a separate Anomaly Detection System (ADS) to identify adversarial attacks.
@@
  • 1.
    We propose GAN-Stability, a framework for training stability prediction systems using exclusively stable instances while maintaining high accuracy. To the best of our knowledge, we are the first to develop such a model in the smart grid context.
  • 2.
    We enhanced the capabilities of our model through adversarial training, enabling GAN-Stability to classify adversarial attacks as unstable samples without the need for an external ADS.
  • 3.
    We evaluate our system on a widely used dataset, achieving an accuracy of up to 0.981 in stability prediction . Additionally, the integration of adversarial training allows the model to detect state-of-the-art adversarial attacks with an accuracy of up to 0.989.
  • 4.
    We compare our solution with state-of-the-art models that require both stable and unstable data for training. Our model demonstrated superior or comparable performances, despite the reduced data requirement for training (i.e., only stable data).
  • 5.
    We test GAN-Stability on an affordable microcomputer, showing reasonable training time requirements and an average response time of less than 7ms during testing.
  • 6.
    We make the code of our systems, attacks, and the dataset available at: https://github.com/emadef1/GAN-Stability
Organization. The rest of the paper is organized as follows: Section 2 reviews existing stability prediction systems and their associated security implications. Section 3 introduces the system and threat model, reflecting real-world scenarios encountered during training and potential attacks. Section 4 discusses the adversarial attack methodologies employed in this study. Section 5 details the proposed stability prediction system and its methodology, while Section 6 evaluates our proposed system using different criteria. Section 7 outlines the limitations and discusses potential areas for improvement. Finally, Section 8 concludes the study with some final remarks.
2 Related Works
In this section, we dig into the existing literature concerning stability prediction systems and their security implications. Specifically, we scrutinize established methodologies for stability prediction in Section 2.1, while looking at the current landscape of attacks targeting these systems in Section 2.2.
2.1 Smart Grid Stability Prediction using AI
@@ 2.2 ML Adversarial Attacks
The paper by Tian et al. [32] investigates security concerns of neural network-based state estimation in smart grids, focusing on adversarial attacks and proposing an efficient method for executing these attacks. Sayghe et al. [33] investigate the impact of adversarial examples on the detection of False Data Injection Attacks (FDIAs) using DL algorithms. Their research examines the repercussions on MLP when exposed to two different adversarial attack strategies. Ahmadian et al. [34] introduced a FDIA using a GAN framework, where the attacker acts as the generative network and the Energy System Operator (ESO) serves as the discriminative network. The attacker generates deceptive data to evade detection by the power system state estimator through an optimization process. Li et al. [35] show that well-established ML models used in energy theft detection systems are susceptible to adversarial attacks. They develop a method to create adversarial measurements, allowing attackers to report significantly lower power consumption to utility companies and evade detection by ML-based systems. In addition, Song et al. [20] conducted a comprehensive analysis of adversarial example attacks in the context of voltage stability assessment for the New England 10-machine 39-bus system. Their study evaluated the reliability of six key attack methods, revealing that most could reduce the target deep neural network’s accuracy by approximately 50% when modifying only half of the input dimensions.
3 System and Threat Model
@@ 3.1 System Model
In an operational setting without active threats, a DSGC stability prediction system assesses whether the grid remains stable or unstable, particularly in a decentralized smart grid context where electricity prices are tied to grid frequency that carry all necessary information about the current power balance. In fact, the stability of electrical grids depends on the balance between electricity generation and demand [7, 10]. In the context of DSGC, stability is characterized by synchronized node frequencies ($\omega$) and steady power flows ($P_{jk}$) across the grid. Stability requires minimal angular frequency deviations and effective damping to suppress oscillations. It is evaluated by the system’s ability to return to equilibrium after disturbances, as measured by mathematical models such as local stability (i.e., linear stability exploring dynamical stability around the steady-state operation of the grid) [10]. Non-stable behaviors arise when synchronization is lost, resulting in significant frequency deviations, amplified oscillations, and destabilized power flows. These effects can lead to cascading failures, particularly when delays, resonance effects, or insufficient damping prevent the grid from recovering.
@@
Our system utilizes ML and AI algorithms to perform binary classification, categorizing grid samples into stable or unstable classes based on various input data collected by the control center from nodes on the grid. Examples of such data include the reaction time of each participant, which indicates how quickly consumers or systems respond to changes; price elasticity coefficients, which reflect the sensitivity of power consumption to changes in electricity prices; and nominal power consumption and production features, which represent baseline levels of power used or generated by the system. As we will discuss in Section 5.1, our system employs 12 different features, typically collected by low-cost equipment from individual prosumers and sent to the control center. Before deployment, the model is trained on clean, uncorrupted data to ensure reliable predictions. Stable grid instances are easy to obtain from operational data as they represent the majority of the system’s operating time. In contrast, collecting unstable data is more challenging. Instability instances require careful labeling by human experts, and acquiring a comprehensive dataset demands long-term observation and significant resources.
@@
In our system model, we adopt a pragmatic approach where we collect enough instances of stability from low-cost equipment by particular prosumers [7]. These instances, all from the same stable label, serve as the sole data for training the stability prediction model. Therefore, to ensure the model’s effectiveness, we should collect a sufficient amount of data that is both comprehensive and representative of the underlying distribution of stable grid conditions. By focusing solely on stability instances, we streamline the training process and alleviate the need for extensive data collection efforts associated with capturing instances of instability. To the best of our knowledge, all the models in the existing literature typically rely on access to both labels in the dataset to make accurate predictions [1, 3, 14, 36, 37, 38]. Our approach challenges this conventional paradigm by demonstrating that accurate stability prediction can be achieved using only instances of stability for training.
3.2 Threat Model
@@
    White-box Scenario: In this scenario, the attacker possesses comprehensive access to both the data employed in testing the model and detailed information regarding the model’s architecture and parameters. This advantageous position provides the attacker with ample opportunities to exploit vulnerabilities in the system. By leveraging this intelligence, the attacker can meticulously craft powerful adversarial samples aimed at deceiving the model. Such a situation could arise if an attacker compromises the control center of the smart grid, for instance, through malware infiltration targeting electric power systems—such as Industroyer [40], which was used in the Ukrainian power grid attack. Additionally, physical attacks on power companies [41] represent another real-world threat that could enable such an attack.
  • 2.
@@
    Grey-box Scenario 1: In this scenario, the attacker has access to the testing data but lacks access to the architecture and parameters of the main model. Despite this limitation, the attacker can still conduct evasion attacks by employing a surrogate model—an alternative model trained on the same dataset—with different architectures. The effectiveness of these attacks relies on transferability properties or the chosen architecture. This condition can exist if an attacker compromises enough prosumers or entities in the system, thus gaining access to several data points but without knowledge of the actual employed model. IoT botnets are an example of how such scenarios could easily become a reality [42]. By compromising large numbers of devices, IoT botnets can collect data or probe the system, potentially providing the attacker with enough information to train a surrogate model. Additionally, attacks such as the Man-in-the-Middle (MitM) attack have been shown to disrupt communication between control systems and field equipment in smart grids. For instance, [43] demonstrates how MitM attacks can leverage false data injection techniques to alter transmitted data, including issuing deceptive commands to field devices.
@@
    To generate adversarial samples, we utilize the LSTM model proposed in [3] as a surrogate model. These adversarial samples will then be deployed against our primary stability prediction system (GAN-Stability). This setup simulates a real-world scenario where potential attackers have access to limited information about the system.
  • 3.
    Grey-box Scenario 2: In this scenario, proposed in [3] for the GAN-GRID attack, the adversary does not have access to real data or the model architecture, but can query the model. The generator neural network is trained using reinforcement learning to produce data that the model classifies as stable. An example of this attack could occur by compromising a grid operator’s dashboard through an insider attack [44], or leveraging malware like CRASHOVERRIDE [45], creating a backdoor for unauthorized access to these systems.
4 Reference Attacks